OpenStudy (anonymous):

This statement tan theta= -12/5, cosecant theta=-13/5 the terminal point is in quadrant 3

OpenStudy (anonymous):

@DragonFlame

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

:/ I don't know how to do this kind of question, but maybe I can tag someone who can hang on

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Zale101

OpenStudy (unklerhaukus):

@wio

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Remember that \(\tan\) is the opposite over adjacent, so you want to draw a triangle with opposite side as -12 and adjacent side as 5.

OpenStudy (anonymous):



OpenStudy (anonymous):

cosecant is going to be hypotenuse over opposite

OpenStudy (anonymous):

Thanks let me calculate really quick

OpenStudy (anonymous):

\[\cos^{-1} \theta =\left(\begin{matrix}x \\ 5\end{matrix}\right)\] Right?

OpenStudy (anonymous):

@wio